the length of a rectangle is 2 inches more that its width. the area of the rectangle is 15 square inches. what are the length and width of the rectangle
Answer by Greenfinch(383)

Width is x, length is x+2, Area is x(x+2) = 15
(x + 5)(x - 3) = 0
Width is 3 inches, Length is 5 inches